# Characterising 1-Rectifiable metric spaces via Connected Tangent Spaces

### v1

@@ Abstract.
We prove that in a complete metric space $X$, $1$-rectifiability of a set $E\subset X$ with $\mathcal{H}^{1}(E)<\infty$ and positive lower density $\mathcal{H}^{1}$-a.e. is implied by the property that all tangent spaces are connected metric spaces.
@@ 1. Introduction
Let $\mathcal{C}_{*}$ be the space of pointed metric measure spaces $(X,d,\mu,x)$ for which the support of $\mu$ is connected.
@@
The fact that 1) implies 2) is given in [2, Theorem 6.6] by applying the results of Kirchheim [5]. Since 2) implies 3) is immediate, our attention is focused on the implication 3) implies 1). We emphasise that, even in the case that all tangent spaces are supported on $\mathbb{R}$, our proof follows a completely different approach that of [2].
Indeed, the geometric idea of our proof begins with a result o Besicovitch [3, Lemma 9] that covers a purely 1-unrectifiable subset of $\mathbb{R}^{2}$ by a large number of circle pairs. A circle pair of a set $F$ is simply $B(x,|x-y|)\cup B(y,|x-y|)$, where $x$ and $y$ are points in $F$. Besicovitch demonstrates that for a $1$-unrectifiable set $F\subset\mathbb{R}^{2}$ and any $r>0$, one may find disjoint circle pairs, whose radii are at most $r$ and sum to at least $\mathcal{H}^{1}(F)/9$, such that the common regions are disjoint from $F$:
@@
To conclude the introduction we mention the related work of Preiss and Tišer [12] that gives a different adaptation of Besicovitch’s circle pair argument (to metric spaces) in order to make progress on the Besicovitch one-half conjecture. These results, and the progress towards the conjecture, were recently extended further by De Lellis, Glaudo, Massaccesi and Vittone [6].
@@ 2.2. Rectifiability
For any $s$, Hausdorff measure is a Borel regular measure [10, Corollary 4.5].
@@
$\mathcal{H}^{s}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})+\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})$ $=\mathcal{H}^{s}(E)$
@@
$=\mathcal{H}^{s}_{\delta}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})+\mathcal{H}^{s}_{\delta}(E\cap\bigcup_{i\in\mathbb{N}}A_{i})+\varepsilon$
$\leq\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})+\sum_{i}\textnormal{diam}(A_{i})^{s}+\varepsilon.$
Subtracting $\mathcal{H}^{s}(E\backslash\bigcup_{i\in\mathbb{N}}A_{i})$ from both sides (which is finite by assumption) gives the desired result. ∎
@@
Given a metric space $(X,d)$ and a set $A\subset\mathbb{R}^{n}$, a map $f:A\rightarrow X$ is Lipschitz if there exists $K>0$ such that
@@
The proof for $X=\mathbb{R}^{m}$ is given in [10, Theorem 15.6]. The proof for general $X$ is identical. ∎
@@ 2.3. Density and Doubling
    If $E$ is $\mathcal{H}^{s}$-measurable, then $\Theta^{*s}(E,x)=0$ for $\mathcal{H}^{s}$-almost all $x\in X\backslash E.$
@@ 2.4. Pointed Gromov-Hausdorff distance
A pointed metric space $(X,\,d,\,x)$ is an ordered triple consisting of a metric space $(X,\,d)$ and a point $x\in X$. A pointed metric measure space $(X,\,d,\,\mu,\,x)$ consists of a complete and separable pointed metric space $(X,x)$ and a measure $\mu\in\mathcal{M}_{loc}(X)$.
An isometric embedding $(X,\,d,\,x)\rightarrow(Y,\,\rho,\,y)$ of pointed metric spaces is an isometric embedding $(X,\,d)\rightarrow(Y,\,\rho)$ such that $x=y.$
@@
[…]
@@
Notation 2.16.
In general we shal not distinguish between the original objects $X$ and $\mu$ and their isometric images $\iota(X)\subset Z$ and $\iota_{\#}\mu$.
@@
Let $(X,\,x)$ and $(Y,\,y)$ be pointed metric spaces. We define the pointed Gromov-Hausdorff distance as
@@
If $\mathbb{M}_{p}$ is the set of isometry classes of proper pointed metric spaces, then $(\mathbb{M}_{p},\,d_{pGH})$ is a complete and seperable metric space. In particular, if $(X,\,x)$ and $(Y,\,y)$ are proper with
@@ 2.5. Convergence of measures
This is a complete and separable metric on the set of isometry classes of pointed metric measure spaces $\mathbb{M}_{pm}$, see [2, Corollary 4.6]. We are now ready to introduce the metric that will be used to define tangent measures in the next section.
@@
Then as shown in [2, Corollary 4.13], $(\mathbb{M}_{*},d_{*})$ is a complete and separable metric space. In light of this, to simplify notation we will sometimes write $(\mu,x)$ in place of $(X,\mu,x)$.
@@
    $\max\{\mu(B(x,1/\varepsilon)\backslash K_{\mu}),\nu(B(y,1/\varepsilon)\backslash K_{\nu})\}<\varepsilon$;
@@ 2.6. Tangents of metric measure spaces
Tangent measures were first introduced by Preiss [11] in Euclidean space as a natural setting in which to explore questions around density. We refer the reader to [10, Page 184] for an introduction to tangent measures. With the notion of convergence of metric measure spaces established in the previous section, the definition of tangent measures in metric spaces follows analogously. First, let $(\mu,d,x)\in\mathbb{M}_{*}$ with $x\in\operatorname{supp}\mu$ and for $r>0$, define
@@
    $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)\quad\implies\quad(C\nu,s\rho,y)\in\textnormal{Tan}(\mu,x)$
@@
Proof of (2) may be found in [2, Proposition 5.5] so we just present a proof of (1). If $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)$ then there exists $r_{i}\rightarrow 0$ such that
@@ 3. Connected Tangents of Metric Spaces
$A=\{x\in\operatorname{supp}\mu\cap B(a,Rd(a,b)):\\ \textnormal{there exists a $\delta$-quasi-path between $a$ and $x$ in $B(a,Rd(a,b))$}\}$
@@
Then for any $K_{\mu}\subset X$ containing $x,\,y$ and satisfying
@@
Let $(X,\,d)$ be a complete metric space and $\mu\in\mathcal{M}_{loc}(X)$. Suppose that for $\mu$-a.e. $x\in\operatorname{supp}\mu$ we have
@@
We use the $(\nu_{n},a_{n})$ to construct another sequence of elements of $\textnormal{Tan}(\mu,x)$.
By Lemma 2.31(1), we may rescale the metrics of each tangent and suppose that $\rho_{n}(a_{n},b_{n})=1$ (we omit the subsequent scaling of the measure by $C$ for ease of notation). By compactness of $\textnormal{Tan}(\mu,x)$, there exists a subsequence $(\nu_{n_{k}},\rho_{n_{k}},a_{n_{k}})$ such that, for any $\varepsilon>0$, there exists a $K>0$ such that
@@
Let $y\in\operatorname{supp}\mu$. We will demonstrate that there exists a $\delta$-quasi-path between $x$ and $y$.
First, by Proposition 2.21, we know that $\mu\in\mathcal{A}(C)$. By Proposition 2.20, $y\in\operatorname{supp}\mu$ implies that there exists a sequence $y_{i}\in\operatorname{supp}\mu_{i}$ such that $y_{i}\rightarrow y$. Since for each $i$ we have that $\mu_{i}\in\mathcal{Q}_{*}(\delta,R,C)$, we also have a sequence of $\delta$-quasi-paths $\{y_{i}^{n}\}_{n\in\mathbb{N}}\subset\operatorname{supp}\mu_{i}\cap B(x,Rd(x,y_{i}))$.
@@
which directly contradicts item (1), concluding the proof.
@@
$F_{R}=\{x\in E:\textnormal{Tan}(\mathcal{H}^{1}\llcorner E,x)\subset\mathcal{Q}_{*}(\delta,R,1/\hat{d})\}.$
@@
which gives us S2. as required.
@@ 4. Besicovitch Partitions
In place of the Besicovitch covering theorem, we use the Alhfors type estimate (4.1) to bound the size of the cover.
@@
For a collection $\mathcal{A}$ and a set $A\in\mathcal{A}$, let
$A_{\cap}:=\{B\in\mathcal{A}:A\cap B\neq\emptyset\}.$
@@
Let $E\subset X$ be a compact set such that for some $r_{0},C>0$ we have
@@
Then for any $r<r_{0}$, there exists a cover $\mathcal{B}$ of $E$ by balls such that
@@
We now consider continua in metric space. Recall that a continuum $L\subset X$ is a compact connected set with $\mathcal{H}^{1}(L)<\infty$.
@@
By [1, Theorems 4.2.1 & 4.4.8], the $L^{i}$ may be covered by 1-Lipschitz images of $[0,2l]$. By [1, Theorem 4.4.3], there exists a subsequence of these maps converging uniformly to some $\gamma\colon[0,2l]\to F$. Setting $L=\gamma([0,l])$ gives the desired result. ∎
The following lemma is a variant of [3, Lemma 8]. In place of an ambient Euclidean space, we will use the Kuratowski embedding to isometrically embed a separable metric space $X$ into $\ell_{\infty}$, the space of real bounded sequences with the supremum norm. In order to use compactness arguments, we note that, if $F\subset\ell_{\infty}$ is compact, then so is $CC(F)$, the closed convex hull of $F$.
@@
We may additionally assume that $\textnormal{dist}(CC(F),x)\leq 1/i$ for all $x\in L^{i}$. Indeed, since each $L_{\boldsymbol{\underline{\nicefrac{{1}}{{i}}}}}^{i}$ is convex, we may replace any part of $L^{i}$ that lies outside by the line segment connecting the end points of this part, creating a shorter curve. Thus we may apply Lemma 4.3 to find that there exists a continuum $L\subset CC(F)$, such that $\mathcal{H}^{1}(L)\leq l$ and such that (for a relabelled subsequence of $L^{i}$ and) all $m\in\mathbb{N}$, there exists an $I>0$ such that
@@
A Besicovitch partition of a set $E\subset X$ is a collection $P=(E_{1},E_{2},\pi,p_{1},p_{2})$ containing
@@
    A distance: $0<\pi=\textnormal{dist}(E_{1},E_{2})$;
@@
    A pair of points: $p_{1}\in E_{1}$ and $p_{2}\in E_{2}$ such that $d(p_{1},p_{2})=\pi$.
We require a notion of disjointness of Besicovitch partitions at a scale corresponding to the $\delta$-path radius that we found via Lemma 3.7. In this vein, we say that two Besicovitch partitions $P=(E_{1},E_{2},\pi,p_{1},p_{2})$ and $P^{\prime}=(E^{\prime}_{1},E^{\prime}_{2},\pi^{\prime},p^{\prime}_{1},p^{\prime}_{2})$ are $R$-disjoint if $B(p_{i},R\pi)\cap B(p^{\prime}_{j},R\pi^{\prime})=\emptyset$ for $i,j\in\{1,2\}$.
The following result is in the spirit of [3, Lemma 9]. However, the construction of the continua in the proof differs quite significantly from the original, which uses the fact that, in the plane, the boundary of a ball is a continuum.
@@
For any $\varepsilon,\delta>0$ and $R\geq 1$, there exists a finite collection $\mathcal{P}$ of Besicovitch partitions $P_{i}=(F_{i,1}F_{i,2},\pi_{i},p_{i,1},p_{i,2})$, $1\leq i\leq n$, such that
@@
    $\sum_{1\leq i\leq n}\pi_{i}>\frac{1-\varepsilon}{3(2R+1)}\mathcal{H}^{1}(F).$
@@
(4.5) $\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
Finally, we satisfy the assumptions of Lemma 4.2 so way may find a cover of $F$ by collection of balls $\mathcal{A}_{1}$ that satisfies
@@
    Find points $p_{1}\in F_{1}$ and $p_{2}\in F_{2}$ such that $\pi:=d(p_{1},p_{2})=\textnormal{dist}(F_{1},F_{2})$, so that $(F_{1},F_{2},\pi,p_{1},p_{2})$ is a Besicovitch partition that satisfies P2..
@@
    Set $\mathcal{P}_{k}=\mathcal{P}_{k-1}\cup\{(F_{1},F_{2},\pi,p_{1},p_{2})\}$, $\mathcal{B}_{k}=\mathcal{B}_{k-1}\cup\{B\}$ and $\mathcal{A}_{k+1}=\mathcal{A}_{k}\cup B$.
@@
Since the $A_{i}$ are convex, this process will terminate after $K\leq m$ steps. Finally, define $\mathcal{A}$ by removing from $\mathcal{A}_{K}$ those $A_{i}$ with $A_{i}\subset B$ for some $B\in\mathcal{B}_{K}$. Then $\mathcal{A}\cup\mathcal{B}_{K}$ forms a covering of $F$ such that each connected component of $\bigcup(\mathcal{A}\cup\mathcal{B}_{K})$ contains an $x_{i}$.
@@
:=$\mathcal{B}:=\{B_{i}\in\mathcal{B}_{K}:B_{i}\cap B_{j}=\emptyset\ \forall 1\leq j<i\}.$
@@
$\mathcal{P}:=\{(F_{1},F_{2},\pi,p_{1},p_{2})_{k}\in\mathcal{P}_{K}:B_{k}\in\mathcal{B}\}.$
@@
    $B(a_{n},\pi/2)\subset B_{\boldsymbol{\underline{\hat{r}+\pi/2}}}.$
@@
    $\frac{N}{C}\pi/2\leq\sum_{n=1}^{N}\mathcal{H}^{1}(F\cap B(a_{n},\pi/2))$ $\leq\mathcal{H}^{1}(F\cap B_{\boldsymbol{\underline{\hat{r}+\pi/2}}})$
    $\leq C(3(2R+1)\pi/2+\hat{r}+\pi/2).$
    Since, by contruction, $\pi\geq\hat{r}/4$, this implies
    $N\leq C^{2}(3(2R+1)+2\hat{r}/\pi+1)\leq C^{2}(3(2R+1)+9).$
@@
$\mathcal{H}^{1}(\tilde{L})$ $\leq\mathcal{H}^{1}(L)+96C^{3}\mathcal{H}^{1}(F)+72C^{3}\mathcal{H}^{1}(F)+6C^{2}(R+2)(1-\varepsilon)\mathcal{H}^{1}(F)$
@@
$\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$
@@
$\mathcal{H}^{1}(F\cap\bigcup\mathcal{A})\leq\mathcal{H}^{1}(F\cap\tilde{L}_{\textbf{\text@underline{r}}})<\frac{\varepsilon}{2}\mathcal{H}^{1}(F).$

### v2

@@ Abstract.
We prove that in a complete metric space $X$, 11-rectifiability of a set $E\subset X$ with $\mathcal{H}^{1}(E)<\infty$ and positive lower density $\mathcal{H}^{1}$-a.e. is implied by the property that all tangent spaces are connected metric spaces.
@@ 1. Introduction
Let $\mathcal{C}^{*}$ be the space of pointed metric measure spaces $(X,d,\mu,x)$ for which the support of $\mu$ is connected.
@@
The fact that 1) implies 2) is given in [3, Theorem 6.6] by applying the results of Kirchheim [6]. Since 2) implies 3) is immediate, our attention is focused on the implication 3) implies 1). We emphasise that, even in the case that all tangent spaces are supported on $\mathbb{R}$, our proof follows a completely different approach than that of [3].
Indeed, the geometric idea of our proof begins with a result of Besicovitch [4, Lemma 9] that covers a purely 1-unrectifiable subset of $\mathbb{R}^{2}$ by a large number of circle pairs. A circle pair of a set $F$ is simply $B(x,|x-y|)\cup B(y,|x-y|)$, where $x$ and $y$ are points in $F$. Besicovitch demonstrates that for a $1$-unrectifiable set $F\subset\mathbb{R}^{2}$ and any $r>0$, one may find disjoint circle pairs, whose radii are at most $r$ and sum to at least $\mathcal{H}^{1}(F)/9$, such that the common regions are disjoint from $F$:
@@
To conclude the introduction we mention the related work of Preiss and Tišer [13] that gives a different adaptation of Besicovitch’s circle pair argument (to metric spaces) in order to make progress on the Besicovitch one-half conjecture. These results, and the progress towards the conjecture, were recently extended further by De Lellis, Glaudo, Massaccesi and Vittone [7]. Further, our result can be viewed as a substantial strengthening of a theorem of Azzam and Mourgoglou [2], which states that a doubling metric measure space with positive 1-density and connected support is 1-rectifiable.
@@ 2.2. Rectifiability
For any $s$, Hausdorff measure is a Borel regular measure [11, Corollary 4.5].
@@
[…]
@@
[…]
@@
Given a metric space $(X,d)$ and a set $A\subset\mathbb{R}^{n}$, a map $f:A\rightarrow X$ is Lipschitz if there exists $K\geq 0$ such that
@@
The proof for $X=\mathbb{R}^{m}$ is given in [11, Theorem 15.6]. The proof for general $X$ is identical. ∎
@@ 2.3. Density and Doubling
    If $E$ is $\mathcal{H}^{s}$-measurable, then $\Theta^{*s}(E,x)=0$ for $\mathcal{H}^{s}$-almost all $\E.x\in X\backslash$ E.
@@ 2.4. Pointed Gromov-Hausdorff distance
A pointed metric space (X,d,x)(X,\,d,\,x) is an ordered triple consisting of a metric space (X,d)(X,\,d) and a point $x\in X$. A pointed metric measure space $(X,\,d,\,\mu,\,x)$ consists of a complete and separable pointed metric space $(X,x)$ and a measure $\mu\in\mathcal{M}_{loc}(X)$.
An isometric embedding $(X,\,d,\,x)\rightarrow(Y,\,\rho,\,y)$ of pointed metric spaces is an isometric embedding $(X,\,d)\rightarrow(Y,\,\rho)$ such that $x\mapsto y.$
@@
(2.16) $d_{H(z)}(X,\,Y):=\inf\left\{\varepsilon>0:X\cap B(z,\,1/\varepsilon)\subset Y_{\underline{\varepsilon}},\,Y\cap B(z,\,1/\varepsilon)\subset X_{\underline{\varepsilon}}\right\}.$
@@
Notation 2.17.
In general we shall not distinguish between the original objects $X$ and $\mu$ and their isometric images $\iota(X)\subset Z$ and $\iota_{\#}\mu$.
@@
Let (X,x)(X,\,x) and (Y,y)(Y,\,y) be pointed metric spaces. We define the pointed Gromov-Hausdorff distance as
@@
If $\mathbb{M}_{p}$ is the set of isometry classes of proper pointed metric spaces, then $(\mathbb{M}_{p},\,d_{pGH})$ is a complete and seperable metric space. In particular, if (X,x)(X,\,x) and (Y,y)(Y,\,y) are proper with
@@ 2.5. Convergence of measures
This is a complete and separable metric on the set of isometry classes of pointed metric measure spaces $\mathbb{M}_{pm}$, see [3, Corollary 4.6]. We are now ready to introduce the metric that will be used to define tangent measures in the next section.
@@
Then as shown in [3, Corollary 4.13], $(\mathbb{M}_{*},d_{*})$ is a complete and separable metric space. In light of this, to simplify notation we will sometimes write $(\mu,x)$ in place of $(X,\mu,x)$.
@@
    $max\{\mu(B(x,1/\varepsilon)\backslash K_{\mu}),\nu(B(y,1/\varepsilon)\backslash K_{\nu})\}<\varepsilon$;
@@ 2.6. Tangents of metric measure spaces
Tangent measures were first introduced by Preiss [12] in Euclidean space as a natural setting in which to explore questions around density. We refer the reader to [11, Page 184] for an introduction to tangent measures. With the notion of convergence of metric measure spaces established in the previous section, the definition of tangent measures in metric spaces follows analogously. First, let $(\mu,d,x)\in\mathbb{M}_{*}$ with $x\in\operatorname{supp}\mu$ and for $r>0$, define
@@
Note that this definition of a tangent measure allows for different sequences of radii to produce different tangents - in a sense allowing the tangents to “rotate”. A basic example is that of a spiral in the plane that converges very quickly in the radial direction. In this case, the set of tangent measures at the origin consists of all half lines, equipped with Lebesgue measure.
@@
    $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)\quad\implies\quad(C\nu,s\rho,y)\in\textnormal{Tan}(\mu,x).$
@@
Proof of (2) may be found in [3, Proposition 5.5] so we just present a proof of (1). If $(\nu,\rho,y)\in\textnormal{Tan}(\mu,x)$ then there exists $r_{i}\rightarrow 0$ such that
@@ 3. Connected Tangents of Metric Spaces
[…]
@@
Then for any $K_{\mu}\subset X$ containing x,yx,\,y and satisfying
@@
Let (X,d)(X,\,d) be a complete metric space and $\mu\in\mathcal{M}_{loc}(X)$. Suppose that for $\mu$-a.e. $x\in\operatorname{supp}\mu$ we have
@@
We use the $(\nu_{n},a_{n})$ to construct another sequence of elements of $\textnormal{Tan}(\mu,x)$. By Lemma 2.34(1), we may rescale the metrics of each tangent and suppose that $\rho_{n}(a_{n},b_{n})=1$ (we omit the subsequent scaling of the measure by $C$ for ease of notation). By compactness of $\textnormal{Tan}(\mu,x)$, there exists a subsequence $(\nu_{n_{k}},\rho_{n_{k}},a_{n_{k}})$ such that, for any $\varepsilon>0$, there exists a $K>0$ such that
@@
Let $y\in\operatorname{supp}\mu$. We will demonstrate that there exists a $\delta$-quasi-path between $x$ and $y$. First, by Proposition 2.24, we know that $\mu\in\mathcal{A}(C)$. By Proposition 2.21, $y\in\operatorname{supp}\mu$ implies that there exists a sequence $y_{i}\in\operatorname{supp}\mu_{i}$ such that $y_{i}\rightarrow y$. Since for each $i$ we have that $\mu_{i}\in\mathcal{Q}_{*}(\delta,R,C)$, we also have a sequence of $\delta$-quasi-paths $\{y_{i}^{n}\}_{n\in\mathbb{N}}\subset\operatorname{supp}\mu_{i}\cap B(x,Rd(x,y_{i}))$.
@@
which directly contradicts item (1), concluding the proof. ∎
@@
[…]
@@
which gives us S2. as required. ∎
@@ 4. Besicovitch Partitions
In place of the Besicovitch covering theorem, we use the Ahlfors type estimate (4.3) to bound the size of the cover.
@@
For a collection $\mathcal{B}$ and a set $B\in\mathcal{B}$, let
$B_{\cap}:=\{B^{\prime}\in\mathcal{B}:B\cap B^{\prime}\neq\emptyset\}.$
@@
Let $E\subset X$ be a totally bounded set such that for some $r_{0},C>0$ we have
@@
Then for any $r<r_{0}$, there exists a finite cover $\mathcal{B}$ of $E$ by balls such that
@@
We now consider continua in metric space. Recall that a continuum $L\subset X$ is a compact connected set with $\mathcal{H}^{1}(L)<\infty$. Recall also that a set with an underlined subscript denotes a closed neighbourhood that set:
$L_{\boldsymbol{\underline{\varepsilon}}}:=\{x\in X:\textnormal{dist}(x,L)\leq\varepsilon\}.$
@@
By [1, Theorems 4.2.1 & 4.4.8], the $L^{i}$ may be covered by 1-Lipschitz images of $[0,2l]$. By [1, Theorem 4.4.3], there exists a subsequence of these maps converging uniformly to some $\gamma\colon[0,2l]\to F$. Setting $L=\gamma([0,2l])$ gives the desired result. ∎
The following lemma is a variant of [4, Lemma 8]. In place of an ambient Euclidean space, we will use the Kuratowski embedding to isometrically embed a separable metric space $X$ into $\ell_{\infty}$, the space of real bounded sequences with the supremum norm. In order to use compactness arguments, we note that, if $F\subset\ell_{\infty}$ is compact, then so is $CC(F)$, the closed convex hull of $F$.
@@
We may additionally assume that $\textnormal{dist}(CC(F),x)\leq 1/i$ for all $x\in L^{i}$. Indeed, because $CC(F)$ is convex, we may replace any part of $L^{i}$ that lies outside by the line segment connecting the end points of this part, creating a shorter curve. Thus we may apply Lemma 4.4 to find that there exists a continuum $L\subset CC(F)$, such that $\mathcal{H}^{1}(L)\leq l$ and such that (for a relabelled subsequence of $L^{i}$ and) all $m\in\mathbb{N}$, there exists an $I>0$ such that
@@
A Besicovitch partition of a set $E\subset X$ is a collection $P=(E_{1},E_{2},\omega,p_{1},p_{2})$ containing
@@
    A distance: $0<\omega=\textnormal{dist}(E_{1},E_{2})$;
@@
    A pair of points: $p_{1}\in E_{1}$ and $p_{2}\in E_{2}$ such that $d(p_{1},p_{2})=\omega$.
We require a notion of disjointness of Besicovitch partitions at a scale corresponding to the $\delta$-path radius that we found via Lemma 3.12. In this vein, we say that two Besicovitch partitions $P=(E_{1},E_{2},\omega,p_{1},p_{2})$ and $P^{\prime}=(E^{\prime}_{1},E^{\prime}_{2},\omega^{\prime},p^{\prime}_{1},p^{\prime}_{2})$ are $R$-disjoint if $B(p_{i},R\omega)\cap B(p^{\prime}_{j},R\omega^{\prime})=\emptyset$ for $i,j\in\{1,2\}$.
The following result is in the spirit of [4, Lemma 9]. However, the construction of the continua in the proof differs quite significantly from the original, which uses the fact that, in the plane, the boundary of a ball is a continuum.
@@
For any $\varepsilon,\delta>0$ and $R\geq 1$, there exists a finite collection $\mathcal{P}$ of Besicovitch partitions $P_{i}=(F_{i,1}F_{i,2},\omega_{i},p_{i,1},p_{i,2})$, $1\leq i\leq n$, such that
@@
    $\sum_{1\leq i\leq n}\omega_{i}>\frac{1-\varepsilon}{3(2R+1)}\mathcal{H}^{1}(F).$
@@
[…]
Finally, we satisfy the assumptions of Lemma 4.2 so way may find a cover of $F$ by a finite collection of balls $\mathcal{A}_{1}$ that satisfies
@@
    Find points $p_{1}\in F_{1}$ and $p_{2}\in F_{2}$ such that $\omega:=d(p_{1},p_{2})=\textnormal{dist}(F_{1},F_{2})$, so that $(F_{1},F_{2},\omega,p_{1},p_{2})$ is a Besicovitch partition that satisfies P2..
@@
    Set $\mathcal{P}_{k}=\mathcal{P}_{k-1}\cup\{(F_{1},F_{2},\omega,p_{1},p_{2})\}$, $\mathcal{B}_{k}=\mathcal{B}_{k-1}\cup\{B\}$ and $\mathcal{A}_{k+1}=\mathcal{A}_{k}\cup\{B\}$.
@@
Since the original cover $\mathcal{A}_{1}$ is finite, this process will terminate after $K\leq\#\text{Cluster}\,{\mathcal{A}}$ steps. Finally, define $\mathcal{A}$ by removing from $\mathcal{A}_{K}$ those $A_{i}$ with $A_{i}\subset B$ for some $B\in\mathcal{B}_{K}$. Then $\mathcal{A}\cup\mathcal{B}_{K}$ forms a covering of $F$ such that each connected component of $\bigcup(\mathcal{A}\cup\mathcal{B}_{K})$ contains an $x_{i}$.
@@
$\mathcal{B}:=\{B_{i}\in\mathcal{B}_{K}:B_{i}\cap B_{j}=\emptyset\ \forall 1\leq j<i\}.$
@@
$\mathcal{P}:=\{(F_{1},F_{2},\omega,p_{1},p_{2})_{k}\in\mathcal{P}_{K}:B_{k}\in\mathcal{B}\}.$
@@
    $B(a_{n},\omega/2)\subset B_{\boldsymbol{\underline{\hat{r}+\omega/2}}}.$
@@
    $\frac{N}{C}\omega/2\leq\sum_{n=1}^{N}\mathcal{H}^{1}(F\cap B(a_{n},\omega/2))$ $\leq\mathcal{H}^{1}(F\cap B_{\boldsymbol{\underline{\hat{r}+\omega/2}}})$
    $\leq C(3(2R+1)\omega/2+\hat{r}+\omega/2).$
    Since, by contruction, $\omega\geq\hat{r}/4$, this implies
    $N\leq C^{2}(3(2R+1)+2\hat{r}/\omega+1)\leq C^{2}(3(2R+1)+9).$
@@
[…]
@@
[…]
@@
[…]